# Same or Not? Enhancing Visual Perception in Vision-Language Models

### v1

@@ 1 Introduction
Fine-grained visual understanding – recognizing subtle details within images – is a hallmark of human vision and a desirable goal for machine perception. Modern vision-language models (VLMs) have shown tremendous progress in broad visual reasoning, however their perception remains coarse-grained: they exhibit systematic biases [vlind, visioncheckup, vlmbias, bscore] and overlook subtle details [blink, vlmsblind]. Fig. 1(a) showcases these challenges. While the vacuum cleaners share the same color and brand name, they are distinct instances, distinguishable by the geometry of their dustbins, handle, and base. Qwen2.5-VL [qwen2.5vl], a strong open-source VLM, incorrectly identifies them as identical and reveals flawed reasoning about their visual attributes.
We attribute the limitations of current VLMs in fine-grained perception partly to their training data. Most large-scale image–text corpora emphasize general visual reasoning – such as spatial relations [clevr, marsili2025visual, gqa, cambrian1], common knowledge [mmmu, okvqa], grounding [visualgenome, flickr30k, dci, refcoco, vcr], or mathematical reasoning [mathqa, mathvista, mathverse, math-vision-dataset] – over detailed visual discrimination. While these datasets enable broad understanding, they provide little incentive to attend to subtle, instance-level differences. This imbalance is stark in leading open-data VLMs [perceptionlm, smolvlm], where the majority of their training data does not reward visual distinctions beyond category-level understanding.
@@
To evaluate progress in fine-grained visual perception in VLMs, we introduce Fgvqa, a benchmark suite for fine-grained VQA. While fine-grained recognition tasks have long existed in vision research, they are typically formulated as 1-of-N classification or retrieval problems, making them ill-suited for VLMs. Fgvqa adapts and reformulates existing recognition and retrieval datasets for VLM evaluation, comprising 12,00012{,}000 queries from diverse domains: Met [met] with artworks such as paintings and sculptures, Inquire [inquire] featuring animals and plants from iNaturalist [inaturalist], Cub [cub] with bird species, Ilias [ilias] showcasing retail products, and Landmarks [landmarksv2] containing landmarks such as monuments and architectural structures.
@@ 2 Related Work
Limitations of VLMs. VLMs trained on large text-image corpora have shown remarkable capabilities in image understanding. Here, proprietary systems [gpt-5, gpt4, gemini2.5, claude] lead in performance, while open-source alternatives [vila, deepseekvl2, llama3, molmo, moondream2, smolvlm, gemma] are narrowing the gap. Despite these advances, VLMs exhibit systematic biases [vlmbias, vlind, bscore, visioncheckup], struggle with fine-grained visual perception [blink, vlmsblind], and often hallucinate details [pope, eyeswideshut, autohallusion, beaf]. To address these challenges, we introduce a large-scale VQA dataset to enhance fine-grained visual understanding in VLMs along with a benchmark suite to evaluate this capability.
@@
VLM Training Data. While most proprietary VLMs rely on closed-source training data [gpt-5, gpt4, claude, gemini2.5, gemini], recent efforts have prioritized open-data alternatives. Molmo [molmo] introduced an open-data model competitive with proprietary systems, while SmolVLM [smolvlm] and PerceptionLM [perceptionlm] scale this paradigm using diverse datasets spanning captioning [coco, nocaps, young2014image, docci, altogether], grounding [visualgenome, flickr30k, dci, refcoco, vcr], spatial reasoning [blink, clevr, vsr], math [intergps, gsm8k, mathqa], and chart/diagram understanding [chartqa, charttotext, ai2d, vistext, charxiv]. Most relevant to us are SpotTheDiff [spotthediff] and Birds-to-Words [birdstowords], which pair images with text describing differences. In contrast, Twin is more comprehensive and larger in size: Twin features both matching and non-matching pairs, enabling reasoning about sameness and difference, distinctions in Twin are finer and pertain to subtle cues, Twin is far larger (561561K pairs vs. 1313K of SpotTheDiff and 3.33.3K of Birds-to-Words), and it spans diverse object types beyond a single domain such as birds. A detailed comparison is in the Appendix.
@@
Fine-grained Recognition categorizes images into specific subgroups [finegrainedsurvey], with applications in classifying plants [inaturalist, oxfordflowers, inquire], birds [cub2011, nabirds], vehicles [fgvcaircrafts, stanfordcars] and recently, retail products [imaterialist, products10k, aliproducts, ilias]. These tasks are typically framed as 1-of-N classification or retrieval, making them ill-suited for evaluating VLMs. We repurpose popular benchmarks [ilias, met, inquire, cub, landmarksv2] and design a task to evaluate the fine-grained VQA abilities of VLMs.
@@
VLM Post-training. Post-training refines VLMs by instilling specialized capabilities and aligning them more closely with human intent. Supervised fine-tuning (SFT) [star, s1, vocot] and distillation [llava-cot, visprogdistill, ical] transfer task-specific skills from annotated data or teacher models, but often cause forgetting and reduced generalization [rlsrazor, sftmemorizesrlgeneralizes]. Reinforcement learning mitigates these issues, advancing instruction-following [rlhf, dpo, qwen3], reasoning [deepseek-r1, deepseekvl2, drgrpo, qwen3, gemini2.5], and spatial understanding [vigorl, grit, marsili2025no]. Beyond post-training methodology, the choice of data plays a critical role in driving post-training success. We demonstrate that reinforcement learning on our fine-grained Twin dataset increases the perceptual precision of VLMs while preserving their performance on core image–language tasks.
@@ 3.1 Dataset Construction
Task Definition. In our task, fine-grained visual understanding requires distinguishing object instances – recognizing when two images show the same object rather than the same category. We define an instance as a set of images of the same physical object under varied viewpoints, lighting, and backgrounds, consistent with prior work [ilias, met, scalablerecognitionvocabtree, landmarksv2, image-retrieval-deep-local]. From this definition, we design a VQA task where a VLM receives two images and determines if they depict the same instance. Unlike category-level recognition (e.g., cat vs dog), our task demands sensitivity to fine, instance-specific cues – shape, texture, and part geometry – that distinguish similar objects. Fig. 1(a) shows the setup; VQA prompts are in the Appendix.
@@
Figure 3: Fgvqa is a suite of fine-grained VQA benchmarks spanning retail products (Twin, Ilias [ilias]); animals and plants (Inquire [inquire]); Landmarks [landmarksv2]; birds (Cub [cub]); and art (Met [met]). We include two query types: pair (top row), where a VLM judges if two images depict the same instance, species, or landmark, and multi where it counts how many images match a reference.
@@
Sourcing Instances. We source object instances across diverse categories from Amazon Reviews [amazon-reviews]. For each object instance, we collect images with varied contexts, viewpoints, and lighting to ensure visual diversity. Human annotators verify that each image clearly shows and identifies the object. We generate positive pairs by sampling from images of the same instance. Fig. 2(a) shows examples of positive pairs along with the dataset’s distribution.
@@
Hard Negative Pairs. A balanced dataset for fine-grained understanding requires both positive and negative pairs. If all examples were positive, the task would be trivial. Likewise, random negatives are often too easy (e.g., a mug paired with a fan). We therefore focus on hard negatives – distinct objects that appear similar. We collect these hard negatives with the help of human annotators. First, using CLIP [clip], we compute pairwise cosine similarities between image embeddings to shortlist visually similar candidates, from which human annotators select the final pairs. Examples of such hard negative pairs are shown in Fig. 2(a), including two white vases with slightly different shapes and two white chairs differing in texture and armrest design.
@@
Generated Negatives. Collecting hard negatives is costly and difficult to scale, so we augment Twin with synthetic negatives using personalized image generation, DreamBooth [dreambooth]. Given images of an instance, DreamBooth produces similar variants that retain overall appearance but alter fine details, simulating hard negatives without extra data collection. The rightmost speaker in Fig. 2(a) demonstrates this: its color and geometry match the left one, but it omits the red logo and lighter contours. Each synthetic image is verified by human annotators. Details on image generation and additional examples are in the Appendix. We validate the impact of these generated negatives in our experiments.
@@ 3.2 Training VLMs on Twin
To improve fine-grained understanding in VLMs and evaluate the impact of our new dataset, we post-train existing models on Twin. Our task requires recognizing subtle attributes to distinguish similar instances, and we hypothesize that optimizing for this task enhances broader fine-grained abilities. We post-train with RL as it is shown to improve model capabilities while preserving prior skills [sftmemorizesrlgeneralizes, rlsrazor, huan2025does, unveilingcompositiongap]. Fig. 4 outlines our model training approach.
@@
Optimization. We tune our VLM $\pi_{\theta}$ using GRPO [deepseekmath] as it balances maximizing expected advantages while preventing drift from the pre-trained VLM. Details are in the Appendix.
Implementation. We train Qwen2.5-VL-3B-Instruct [qwen2.5vl] and InternVL3.5-1B-Instruct [internvl3] on Twin, chosen as leading open-source VLMs at the 3B and 1B scales. Training is end-to-end on $4$ A100 GPUs for $1$ epoch with a batch size of $480$, group size $5$, and learning rate $10^{-6}$. We use verl [verl] for optimization and balance positive and negative queries in each batch. Hyperparameters are in the Appendix.
@@ 4 The Fgvqa Benchmark Suite
Despite their strong general visual abilities, current VLMs struggle with fine-grained understanding [vlmbias, vlmsblind]. Existing VQA benchmarks emphasize broad image understanding, leaving this capability largely unevaluated. To this end, we introduce Fgvqa, a suite of fine-grained VQA benchmarks that measure how well VLMs identify subtle visual cues. We show examples in Fig. 3 and provide more details below.
Benchmarks. Fine-grained understanding is a general skill – models attuned to subtle differences should generalize across domains. To evaluate this, we repurpose recognition and retrieval datasets, totaling 12,00012{,}000 queries spanning diverse domains: Met [met] (artworks such as paintings and sculptures), Ilias [ilias] (retail products), Landmarks [landmarksv2] (monuments), Cub [cub] (bird species), Inquire [inquire] (animals and plants from iNaturalist [inaturalist]), and Twin-Eval, collected like Twin but with unseen instances. The breadth of Fgvqa enables assessment of cross-domain generalization.
Twin-Eval Ilias [ilias] Landmarks [landmarksv2] Met [met] Cub [cub] Inquire [inquire]
@@
PerceptionLM 3B [perceptionlm] $37.8$ $46.9$ $25.2$ $36.1$ $49.6$ $25.0$ $37.3$ $51.5$ $25.7$ $38.6$ $50.4$ $26.1$ $38.3$ $49.6$ $24.9$ $37.3$ $52.1$ $26.1$ $39.1$
Moondream 2 2B [moondream2] $45.9$ $64.5$ $30.0$ $47.3$ $58.7$ $30.3$ $44.5$ $59.8$ $25.2$ $42.5$ $84.3$ $27.0$ $55.7$ $50.6$ $43.3$ $47.0$ $51.9$ $25.1$ $38.5$
Gemma3 4B [gemma3] $68.7$ $68.1$ $39.5$ $53.8$ $88.4$ $46.5$ $67.5$ $91.2$ $55.2$ $73.2$ $87.9$ $57.5$ $72.7$ $91.2$ $55.2$ $73.2$ $90.0$ $48.7$ $69.4$
SmolVLM2 2.2B [smolvlm] $46.9$ $54.7$ $25.1$ $39.9$ $52.2$ $29.7$ $41.0$ $54.1$ $29.6$ $41.9$ $73.9$ $27.9$ $50.9$ $81.5$ $35.0$ $58.3$ $73.0$ $25.7$ $49.4$
@@ 5 Experiments
(1) Proprietary VLMs: We establish baseline performance on Fgvqa for GPT4o [gpt4] and Gemini 2.5 Flash [gemini2.5].
(2) Open-Source VLMs: We report leading open-source models PerceptionLM-3B [perceptionlm], Moondream2-2B [moondream2], Gemma3-4B [gemma3], and SmolVLM2-2.2B [smolvlm].
(3) Direct Comparisons: We fine-tune Qwen2.5-VL 3B Instruct [qwen2.5vl] and InternVL3.5 1B Instruct [internvl3] to evaluate the value of Twin across architectures and scales. Each base model is compared to its trained variant, denoted “+ Twin”.
@@ 5.1 Main Results
Comparison to open-source VLMs. Among open-source baselines, Gemma3 performs best across benchmarks: $53.8\%$ on Twin-Eval, $69.4\%$ on Inquire, and $72.7\%$ on Ilias. However, as Gemma3 does not support training with Flash Attention [flashattention2, gemma-flash-attn1, gemma-flash-attn2], fine-tuning it in PyTorch would be prohibitively expensive. Therefore, we train the second-best open-source VLM, Qwen2.5-VL, on Twin. Our trained Qwen2.5-VL model (+ Twin) outperforms Gemma3 on Twin-Eval ($67.3\%$ vs $53.8\%$), Inquire ($73.7\%$ vs $69.4\%$), and Cub ($75.1\%$ vs $73.2\%$). Gemma3 outperforms our trained Qwen2.5-VL model on Met ($72.7\%$ vs $66.0\%$) and Landmarks ($73.2\%$ vs $57.9\%$) – the datasets where we see the least improvement from training.
@@ 5.2 Ablations and Analysis
SFT vs. RL. We compare post-training with reinforcement learning (RL) versus supervised fine-tuning (SFT) in Tab. 3. Our model produces both a chain-of-thought (CoT) explanation and a final yes/no prediction. RL optimizes a reward based solely on the final answer correctness, while SFT performs supervision on all output tokens, including both CoT and final prediction. To obtain high-quality supervision for SFT, we collect responses from Gemini-2.5-Flash [gemini2.5] on a subset of Twin and retain only those where Gemini is correct. We then post-train Qwen2.5-VL on these 136136K pairs with both SFT and RL. Details for SFT are in the Appendix. Although both methods achieve similar performance on Twin-Eval ($61.2\%$ with SFT vs $64.0\%$ with RL), the RL variant outperforms on out-of-distribution datasets ($53.9\%$ with SFT vs $65.0\%$ with RL on Cub; $52.3\%$ with SFT vs. $58.7\%$ with RL on Met). These results corroborate prior findings that RL better preserves generalization during post-training than SFT [sftmemorizesrlgeneralizes, rlsrazor, huan2025does, unveilingcompositiongap].
Realworld [realworldqa] Pope [pope] Cv-Bench [cambrian1] Nlvr2 [nlvr2] Ai2d [ai2d] Seed [seedbench] ChartQA [chartqa] TextVQA [textvqa] Blink [blink] Mmmu [mmmu] VlmBias [vlmbias]
@@
Pets [oxfordpets] Sun397 [sun397] Cub [cub] Cifar100 [cifar100]
@@
Encoder Probing. We probe the underlying vision encoder of Qwen2.5-VL to examine whether training on Twin improves its visual representations. Unlike contrastive models (e.g. CLIP [clip]), this encoder was not trained to produce embeddings useful in isolation. Nevertheless, we apply traditional probing methods to assess representation quality and determine if post-training on Twin improves representation ability of vision encoders in VLMs. In Tab. 5, we report accuracy on fine-grained classification datasets covering various domains: Pets [oxfordpets] with cat and dog species, Sun397 [sun397] with real world scenes, Cub [cub] with bird species, and Cifar100 [cifar100] with everyday objects and animals. We follow [perceptionencoder] and evaluate two settings: a linear probe (Lin.) trained on embeddings from each dataset’s training split and then frozen for validation, and K-nearest neighbor retrieval (KNN), where test queries are matched against embeddings from the corresponding training split.
From Tab. 5, we find that post-training on Twin yields consistent improvements across fine-grained classification datasets: $4.1\%$ with a linear probe on Pets ($75.0\%$ to $79.1\%$), $1.4\%$ with KNN on Cifar100 ($71.8\%$ to $73.2\%$), and $2\%$ with a linear probe on Sun397 ($64.4\%$ to $66.4\%$). The improvement to probing accuracy across datasets suggest that by forcing VLMs to detect subtle differences in objects, post-training on Twin yields embeddings better suited for fine-grianed understanding.

### v2

@@ 1 Introduction
Fine-grained visual understanding – recognizing subtle details within images – is a hallmark of human vision and a desirable goal for machine perception. Modern vision-language models (VLMs) have shown tremendous progress in broad visual reasoning, however their perception remains coarse-grained: they exhibit systematic biases [38, 74, 89, 90] and overlook subtle details [19, 69]. Fig. 1(a) showcases these challenges. While the vacuum cleaners share the same color and brand name, they are distinct instances, distinguishable by the geometry of their dustbins, handle, and base. Qwen2.5-VL [5], a strong open-source VLM, incorrectly identifies them as identical and reveals flawed reasoning about their visual attributes.
We attribute the limitations of current VLMs in fine-grained perception partly to their training data. Most large-scale image–text corpora emphasize general visual reasoning – such as spatial relations [30, 52, 27, 83], common knowledge [108, 51], grounding [36, 66, 85, 31, 110], or mathematical reasoning [3, 47, 112, 93] – over detailed visual discrimination. While these datasets enable broad understanding, they provide little incentive to attend to subtle, instance-level differences. This imbalance is stark in leading open-data VLMs [10, 50], where the majority of their training data does not reward visual distinctions beyond category-level understanding.
@@
To evaluate progress in fine-grained visual perception in VLMs, we introduce Fgvqa, a benchmark suite for fine-grained VQA. While fine-grained recognition tasks have long existed in vision research, they are typically formulated as 1-of-N classification or retrieval problems, making them ill-suited for VLMs. Fgvqa adapts and reformulates existing recognition and retrieval datasets for VLM evaluation, comprising 12,00012{,}000 queries from diverse domains: Met [107] with artworks such as paintings and sculptures, Inquire [88] featuring animals and plants from iNaturalist [87], Cub [91] with bird species, Ilias [33] showcasing retail products, and Landmarks [97] containing landmarks such as monuments and architectural structures.
@@ 2 Related Work
Limitations of VLMs. VLMs trained on large text-image corpora have shown remarkable capabilities in image understanding. Here, proprietary systems [63, 1, 13, 4] lead in performance, while open-source alternatives [43, 99, 16, 15, 55, 50, 82] are narrowing the gap. Despite these advances, VLMs exhibit systematic biases [89, 38, 90, 74], struggle with fine-grained visual perception [19, 69], and often hallucinate details [41, 84, 98, 105]. To address these challenges, we introduce a large-scale VQA dataset to enhance fine-grained visual understanding in VLMs along with a benchmark suite to evaluate this capability.
@@
VLM Training Data. While most proprietary VLMs rely on closed-source training data [63, 1, 4, 13, 80], recent efforts have prioritized open-data alternatives. Molmo [15] introduced an open-data model competitive with proprietary systems, while SmolVLM [50] and PerceptionLM [10] scale this paradigm using diverse datasets spanning captioning [44, 2, 106, 62, 103], grounding [36, 66, 85, 31, 110], spatial reasoning [19, 30, 45], math [48, 12, 3], and chart/diagram understanding [54, 61, 32, 79, 95]. Most relevant to us are SpotTheDiff [29] and Birds-to-Words [18], which pair images with text describing differences. In contrast, Twin is more comprehensive and larger in size: Twin features both matching and non-matching pairs, enabling reasoning about sameness and difference, distinctions in Twin are finer and pertain to subtle cues, Twin is far larger (561561K pairs vs. 1313K of SpotTheDiff and 3.33.3K of Birds-to-Words), and it spans diverse object types beyond a single domain such as birds. A detailed comparison is in the Appendix.
@@
Fine-grained Recognition categorizes images into specific subgroups [96], with applications in classifying plants [87, 58, 88], birds [92, 86], vehicles [49, 35] and recently, retail products [22, 6, 9, 33]. These tasks are typically framed as 1-of-N classification or retrieval, making them ill-suited for evaluating VLMs. We repurpose popular benchmarks [33, 107, 88, 91, 97] and design a task to evaluate the fine-grained VQA abilities of VLMs.
@@
VLM Post-training. Post-training refines VLMs by instilling specialized capabilities and aligning them with human intent. Supervised fine-tuning (SFT) [109, 56, 42] and distillation [102, 25, 71] transfer task-specific skills from annotated data or teacher models, but often cause forgetting and reduced generalization [75, 11]. Reinforcement learning mitigates these issues, advancing instruction-following [64, 68, 104], reasoning [21, 99, 46, 104, 13], and spatial understanding [72, 17, 53]. Beyond post-training methodology, the choice of data plays a critical role in driving post-training success. We demonstrate that reinforcement learning on our fine-grained Twin dataset increases the perceptual precision of VLMs while preserving their performance on core image–language tasks.
@@ 3.1 Dataset Construction
Task Definition. In our task, fine-grained visual understanding requires distinguishing object instances – recognizing when two images show the same object rather than the same category. We define an instance as a set of images of the same physical object under varied viewpoints, lighting, and backgrounds, consistent with prior work [33, 107, 59, 97, 60]. From this definition, we design a VQA task where a VLM receives two images and determines if they depict the same instance. Unlike category-level recognition (e.g., cat vs dog), our task demands sensitivity to fine, instance-specific cues – shape, texture, and part geometry – that distinguish similar objects. Fig. 1(a) shows the setup; VQA prompts are in the Appendix.
@@
Figure 3: Fgvqa is a suite of fine-grained VQA benchmarks spanning retail products (Twin, Ilias [33]); animals and plants (Inquire [88]); Landmarks [97]; birds (Cub [91]); and art (Met [107]). We include two query types: pair (top row), where a VLM judges if two images depict the same instance, species, or landmark, and multi where it counts how many images match a reference.
@@
Sourcing Instances. We source object instances across diverse categories from Amazon Reviews [24]. For each object instance, we collect images with varied contexts, viewpoints, and lighting to ensure visual diversity. Human annotators verify that each image clearly shows and identifies the object. We generate positive pairs by sampling from images of the same instance. Fig. 2(a) shows examples of positive pairs along with the dataset’s distribution.
@@
Hard Negative Pairs. A balanced dataset for fine-grained understanding requires both positive and negative pairs. If all examples were positive, the task would be trivial. Likewise, random negatives are often too easy (e.g., a mug paired with a fan). We therefore focus on hard negatives – distinct objects that appear similar. We collect these hard negatives with the help of human annotators. First, using CLIP [67], we compute pairwise cosine similarities between image embeddings to shortlist visually similar candidates, from which human annotators select the final pairs. Examples of such hard negative pairs are shown in Fig. 2(a), including two white vases with slightly different shapes and two white chairs differing in texture and armrest design.
@@
Generated Negatives. Collecting hard negatives is costly and difficult to scale, so we augment Twin with synthetic negatives using personalized image generation, DreamBooth [70]. Given images of an instance, DreamBooth produces similar variants that retain overall appearance but alter fine details, simulating hard negatives without extra data collection. The rightmost speaker in Fig. 2(a) demonstrates this: its color and geometry match the left one, but it omits the red logo and lighter contours. Each synthetic image is verified by human annotators. Details on image generation and additional examples are in the Appendix. We validate the impact of these generated negatives in our experiments.
@@ 3.2 Training VLMs on Twin
To improve fine-grained understanding in VLMs and evaluate the impact of our new dataset, we post-train existing models on Twin. Our task requires recognizing subtle attributes to distinguish similar instances, and we hypothesize that optimizing for this task enhances broader fine-grained abilities. We post-train with RL as it is shown to improve model capabilities while preserving prior skills [11, 75, 26, 40]. Fig. 4 outlines our model training approach.
@@
Optimization. We tune our VLM $\pi_{\theta}$ using GRPO [73] as it balances maximizing expected advantages while preventing drift from the pre-trained VLM. Details are in the Appendix.
Implementation. We train Qwen2.5-VL-3B-Instruct [5] and InternVL3.5-1B-Instruct [94] on Twin, chosen as leading open-source VLMs at the 3B and 1B scales. Training is end-to-end on $4$ A100 GPUs for $1$ epoch with a batch size of $480$, group size $5$, and learning rate $10^{-6}$. We use verl [76] for optimization and balance positive and negative queries in each batch. Hyperparameters are in the Appendix.
@@ 4 The Fgvqa Benchmark Suite
Despite their strong general visual abilities, current VLMs struggle with fine-grained understanding [89, 69]. Existing VQA benchmarks emphasize broad image understanding, leaving this capability largely unevaluated. To this end, we introduce Fgvqa, a suite of fine-grained VQA benchmarks that measure how well VLMs identify subtle visual cues. We show examples in Fig. 3 and provide more details below.
Benchmarks. Fine-grained understanding is a general skill – models attuned to subtle differences should generalize across domains. To evaluate this, we repurpose recognition and retrieval datasets, totaling 12,00012{,}000 queries spanning diverse domains: Met [107] (artworks such as paintings and sculptures), Ilias [33] (retail products), Landmarks [97] (monuments), Cub [91] (bird species), Inquire [88] (animals and plants from iNaturalist [87]), and Twin-Eval, collected like Twin but with unseen instances. The breadth of Fgvqa enables assessment of cross-domain generalization.
Twin-Eval Ilias [33] Landmarks [97] Met [107] Cub [91] Inquire [88]
@@
PerceptionLM 3B [10] $37.8$ $46.9$ $25.2$ $36.1$ $49.6$ $25.0$ $37.3$ $51.5$ $25.7$ $38.6$ $50.4$ $26.1$ $38.3$ $49.6$ $24.9$ $37.3$ $52.1$ $26.1$ $39.1$
Moondream 2 2B [55] $45.9$ $64.5$ $30.0$ $47.3$ $58.7$ $30.3$ $44.5$ $59.8$ $25.2$ $42.5$ $84.3$ $27.0$ $55.7$ $50.6$ $43.3$ $47.0$ $51.9$ $25.1$ $38.5$
Gemma3 4B [81] $68.7$ $68.1$ $39.5$ $53.8$ $88.4$ $46.5$ $67.5$ $91.2$ $55.2$ $73.2$ $87.9$ $57.5$ $72.7$ $91.2$ $55.2$ $73.2$ $90.0$ $48.7$ $69.4$
SmolVLM2 2.2B [50] $46.9$ $54.7$ $25.1$ $39.9$ $52.2$ $29.7$ $41.0$ $54.1$ $29.6$ $41.9$ $73.9$ $27.9$ $50.9$ $81.5$ $35.0$ $58.3$ $73.0$ $25.7$ $49.4$
@@ 5 Experiments
(1) Proprietary VLMs: We establish baseline performance on Fgvqa for GPT4o [1] and Gemini 2.5 Flash [13].
(2) Open-Source VLMs: We report leading open-source models PerceptionLM-3B [10], Moondream2-2B [55], Gemma3-4B [81], and SmolVLM2-2.2B [50].
(3) Direct Comparisons: We fine-tune Qwen2.5-VL 3B Instruct [5] and InternVL3.5 1B Instruct [94] to evaluate the value of Twin across architectures and scales. Each base model is compared to its trained variant, denoted “+ Twin”.
@@ 5.1 Main Results
Comparison to open-source VLMs. Among open-source baselines, Gemma3 performs best across benchmarks: $53.8\%$ on Twin-Eval, $69.4\%$ on Inquire, and $72.7\%$ on Ilias. However, as Gemma3 does not support training with Flash Attention [14, 28, 20], fine-tuning it in PyTorch would be prohibitively expensive. Therefore, we train the second-best open-source VLM, Qwen2.5-VL, on Twin. Our trained Qwen2.5-VL model (+ Twin) outperforms Gemma3 on Twin-Eval ($67.3\%$ vs $53.8\%$), Inquire ($73.7\%$ vs $69.4\%$), and Cub ($75.1\%$ vs $73.2\%$). Gemma3 outperforms our trained Qwen2.5-VL model on Met ($72.7\%$ vs $66.0\%$) and Landmarks ($73.2\%$ vs $57.9\%$) – the datasets where we see the least improvement from training.
@@ 5.2 Ablations and Analysis
SFT vs. RL. We compare post-training with reinforcement learning (RL) versus supervised fine-tuning (SFT) in Tab. 3. Our model produces both a chain-of-thought (CoT) explanation and a final yes/no prediction. RL optimizes a reward based solely on the final answer correctness, while SFT performs supervision on all output tokens, including both CoT and final prediction. To obtain high-quality supervision for SFT, we collect responses from Gemini-2.5-Flash [13] on a subset of Twin and retain only those where Gemini is correct. We then post-train Qwen2.5-VL on these 136136K pairs with both SFT and RL. Details for SFT are in the Appendix. Although both methods achieve similar performance on Twin-Eval ($61.2\%$ with SFT vs $64.0\%$ with RL), the RL variant outperforms on out-of-distribution datasets ($53.9\%$ with SFT vs $65.0\%$ with RL on Cub; $52.3\%$ with SFT vs. $58.7\%$ with RL on Met). These results corroborate prior findings that RL better preserves generalization during post-training than SFT [11, 75, 26, 40].
Realworld [100] Pope [41] Cv-Bench [83] Nlvr2 [78] Ai2d [32] Seed [39] ChartQA [54] TextVQA [77] Blink [19] Mmmu [108] VlmBias [89]
@@
Pets [65] Sun397 [101] Cub [91] Cifar100 [37]
@@
Encoder Probing. We probe the underlying vision encoder of Qwen2.5-VL to examine whether training on Twin improves its visual representations. Unlike contrastive models (e.g. CLIP [67]), this encoder was not trained to produce embeddings useful in isolation. Nevertheless, we apply traditional probing methods to assess representation quality and determine if post-training on Twin improves representation ability of vision encoders in VLMs. In Tab. 5, we report accuracy on fine-grained classification datasets covering various domains: Pets [65] with cat and dog species, Sun397 [101] with real world scenes, Cub [91] with bird species, and Cifar100 [37] with everyday objects and animals. We follow [7] and evaluate two settings: a linear probe (Lin.) trained on embeddings from each dataset’s training split and then frozen for validation, and K-nearest neighbor retrieval (KNN), where test queries are matched against embeddings from the corresponding training split.
From Tab. 5, we find that post-training on Twin yields consistent improvements across fine-grained classification datasets: $4.1\%$ with a linear probe on Pets ($75.0\%$ to $79.1\%$), $1.4\%$ with KNN on Cifar100 ($71.8\%$ to $73.2\%$), and $2\%$ with a linear probe on Sun397 ($64.4\%$ to $66.4\%$). The improvement to probing accuracy across datasets suggest that by forcing VLMs to detect subtle differences in objects, post-training on Twin yields embeddings better suited for fine-grained understanding.